OpenStudy (anonymous):

Evaluate the limit. Lim x-->0- (x^2-8)/x??

OpenStudy (zarkon):

\[\lim_{x\to 0^{-}}\frac{x^2-8}{x}=\infty\]

OpenStudy (anonymous):

how do you know?

OpenStudy (anonymous):

like can you explain?

OpenStudy (zarkon):

bottom is getting very small (but is negative)...top is getting close to -8. (so fraction is going to +infty since... neg/neg=pos